OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (anonymous):

except that \[3-x-y\] is not an equation. first you need the slope

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}=\frac{2-1}{5-4}=\frac{1}{1}=1\]

OpenStudy (anonymous):

now use almighty point slope formula \[y=y_1=m(x=x_1)\] in your case you can use \[x_1=4,y_1=1\] to get \[y-1=1(x-4)\] \[y-1=x-4\] \[y=x-3\]

OpenStudy (anonymous):

so it should be 3-x=y